1.True or False: When a ball is twirling at a uniform speed on the end of a string, the inward force is zero.

2.True or False: An object can be accelerating while maintaining a uniform (constant) speed.

3.If you were to reduce the radius of the arc through which you twirled a ball on a string by one-half, the size of the inward force (tension) in the string would be ___ as large

1/4 x

1/2 x

2 x

4 x


4.If you were to increase the mass of the ball by five times, the size of the inward force (tension) in the string would be ___ as large.

1/10 x

1/5 x

5 x

25x

5.The centripetal force acting on a 2.5 kg object whirling at a speed of 3.5 m/s in a horizontal circle of radius 0.80 m is: ___ N. (Give your answer with the correct number of significant digits and do not include units)



1
Expert's answer
2021-07-13T10:42:21-0400

1) False

2) True

3) 2 x

4) 5 x

5)


"F=m\frac{v^2}{r}\\F=2.5\frac{3.5^2}{0.8}=28\ N\\"